OpenStudy (anonymous):

How do I solve this? 3/5f+24=4-1/5f

OpenStudy (anonymous):

multiply throughout by 5f to get 3 + 120f = 96f -1 then rearrange to get: 24f = -4 f = -1/6

OpenStudy (dejavo):

f= -25

OpenStudy (anonymous):

How did u get that

OpenStudy (anonymous):

Can u please show step by step?

OpenStudy (agent0smith):

\[\Large \frac{ 3 }{5f } +24 = 4-\frac{ 1 }{ 5f }\]yes...?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Actually the f is besides the fractions

OpenStudy (agent0smith):

\[\Large \frac{ 3 }{5 }f +24 = 4-\frac{ 1 }{ 5}f\]if you want to get rid of the fractions, multiply both sides by 5 (you must multiply everything by 5)

OpenStudy (anonymous):

Alright

OpenStudy (agent0smith):

So try that, OR just add 1/5f to both sides, and show me what you get